### v1

# Characterizing Continual Learning Scenarios and Strategies for Audio Analysis

Abstract
Audio analysis is useful in many application scenarios. The state-of-the-art audio analysis approaches assume that the data distribution at training and deployment time will be the same. However, due to various real-life environmental factors, the data may encounter drift in its distribution or can encounter new classes in the late future. Thus, a one-time trained model might not perform adequately. In this paper, we characterize continual learning (CL) approaches in audio analysis. In this paper, we characterize continual learning (CL) approaches, intended to tackle catastrophic forgetting arising due to drifts. As there is no CL dataset for audio analysis, we use DCASE 2020 to 2023 datasets to create various CL scenarios for audio-based monitoring tasks. We have investigated the following CL and non-CL approaches: EWC, LwF, SI, GEM, A-GEM, GDumb, Replay, Naive, cumulative, and joint training. The study is very beneficial for researchers and practitioners working in the area of audio analysis for developing adaptive models. We observed that Replay achieved better results than other methods in the DCASE challenge data. It achieved an accuracy of 70.12% for the domain incremental scenario and an accuracy of 96.98% for the class incremental scenario.
Index Terms:
machine monitoring, audio data, data shift, continual learning, lifelong surveillance
I Introduction
Audio plays an important role in machine learning and deep learning applications, contributing to various tasks such as speech recognition [1], audio surveillance [2], sound classification, audio signal processing [3], music generation, and environmental sound analysis. In this paper, we characterize continual learning (CL) approaches for audio analysis with a case study of audio-based monitoring. Audio-based monitoring offers distinct advantages such as low-cost deployment, non-invasiveness, remote operation, and the ability to capture rich contextual information. With advancements in sensor technologies and the widespread availability of audio data, audio-based monitoring has become increasingly prevalent across various domains, including public surveillance, environmental sound monitoring, wildlife conservation, and healthcare.
In the real world, data cannot be obtained at once; it usually comes over time. We need to keep updating the models with new data. However, retraining with new data leads to a phenomenon known as catastrophic forgetting [4][5][6][7]. To mitigate this challenge, we need to use approaches that can remember everything we learned once. Continual learning (CL) is an approach that allows the model to learn new data and remember the previous learning simultaneously [8, 9]. Different CL approaches are used to handle catastrophic forgetting. The approaches compared in this paper are EWC [10], SI [11], LwF [12][13], GEM [14], A-GEM [15], GDumb [16], Replay, Naive, cumulative, and Joint Training. We have used ResNet50 as the backbone model for these approaches.
CL has three fundamental scenarios: task-incremental, domain-incremental, and class-incremental scenario. In this paper, the dataset is arranged in two criteria: domain incremental and class incremental. In domain incremental, the data consists of data samples in different batches or tasks, and each task has similar classes but a different distribution of data samples. While in class incremental, the tasks can have different classes associated with the data distribution. Different CL scenarios are created using data from DCASE [^1] https://dcase.community/ challenge, which provides a huge publicly available audio dataset with annotations.
The major contributions of this paper are as follows:
  • We have given a comparative study of different CL approaches, namely EWC, LwF, SI, GEM, A-GEM, GDumb, Replay, Naive, cumulative, and JointTraining with maximum possible variations of their parameter values to get the best performance of the particular approach for two CL scenarios like domain incremental and class incremental.
  • We have prepared a domain and class incremental scenarios dataset using the publicly available DCASE audio dataset of 2020 to 2023 challenges.
The rest of the paper is structured as follows. Section II contains the literature survey on monitoring and CL approaches, followed by section III, which gives the basics of CL methods and scenarios. Section IV consists of the basics of CL and Non-CL baselines. Section V includes a detailed description of the dataset used. Section VI consists of the experimental setup and results, followed by a discussion in section VII, and the conclusion in section VIII.
II Literature Review
This section provides a brief overview of CL approaches in the context of audio analysis. CL methods enable machine learning models to learn sequentially from a continuous stream of data. CL allows them to acquire new skills and accumulate knowledge over time without the risk of forgetting previously learned experiences, commonly known as catastrophic forgetting [17]. In the audio domain, continual learning becomes pertinent due to the sequential, dynamic, and evolving nature of sound environments, which requires adaptive models. By continuously updating models with new audio samples, CL algorithms can learn to recognize and classify new audio events, adapt to new acoustic conditions, and generalize across different audio contexts. The audio context or applications seen so far are sound classification [18], audio analysis/classification [19] [20][21][22][23], sound event detection [24][25], audio classification [26], audio captioning [27], audio-visual learning [28][29][30], fake audio detection [31], speech recognition [32], etc.
A summary of CL works in the audio domain is provided in Table I. From Table I, we observe that most of the audio-based CL applications focus on audio scene classification, with very few addressing anomaly detection. It can be inferred that the DCASE anomaly dataset utilized in this paper has not been previously employed by others. Although some studies have utilized the DCASE dataset, they have done so for different applications, such as audio scene classification, rather than anomaly detection. For instance, in [33], the DCASE challenge anomaly dataset was utilized for anomaly detection, although in a different manner. They treated subsequent machine types as the drift, whereas we considered naturally occurring drift. Further, we can observe that most works focused on handling the unseen classes or working in class incremental scenarios, and very few handle the domain shift. Moreover, almost all works focus on either scenario, i.e., domain or class incremental scenario, but not both. In this paper, we have studied both scenarios and performed a comparative analysis of various state-of-the-art continual learning strategies.
TABLE I: Summary of state-of-the-art continual learning works in audio domain
Work CL Strategy CL Scenario Dataset
Application (shift detail if any)
Wang et. al.[34]
Rehearsal
CI ESC10 Audio scene classification
Koh et. al.[35] Architectural CI DCASE16, US-SED Audio scene classification
Wang et. al.[36] Architectural CI
FSD-MIX-CLIPS (1s),
FSD-MIX-SED (10s)
Audio scene classification
Kwon et. al.[37] Rehearsal CI
EmotionSense,
UrbanSound8K
Audio scene classification
Wang et. al.[38] Regularization CI
UrbanSound8K, VGGSound,
urban acoustic scenes 2019
(DCASE TAU19)
Audio scene classification
Karam et. al.[39] Rehearsal TI ESC50, UrbanSound8K Audio scene classification
Mulimani et. al.[40] Regularization CI
DCASE TUT acoustic scenes
2017, TUT 2016/2017
Audio scene classification
Sun et. al.[41]
Rehearsal
CI ESC10, ESC50, UrbanSound8K Audio scene classification
Wang et. al.[42] Architectural CI AudioSet, ESC50
Multi-label audio classification
Mulimani et. al.[43] Regularization CI AudioSet
Multi-label audio classification
Mo et. al.[44]
Regularization
& Rehearsal
CI
VGGSound-Instruments,
VGGSound-100,
VGG-Sound Sources
Multi-modal classification (audio,video)
Pian et.al.[45] Rehearsal CI
AVE-CI, K-S-CI,
VS100-CI
Multi-modal classification (audio,video)
Kim et. al.[46] Regularization TI
AudioSet, VGGSound, MACS,
FSD50K, ClothoV2, AudioCaps
Multi-modal classification (audio,video, text)
Berg et. al.[47] Regularization DI Clotho, AudioCaps
Automated audio captioning
(Audio clips are annotated with 5 captions
to each 10-30 sec each of both datasets)
Ma et. al.[48] Regularization DI ASVSpoof2019
Fake audio detection
(Considered 2 subsets of the dataset &
selected 4 spoofing attacks with big
differences, from each subset separately)
Zhou et. al.[33] Regularization DI DCASE 2020 challenge Task2
Anomaly sound detection
(Considered subsequent machine type as drift)
III Continual learning
Continual learning (CL) is a rapidly emerging research direction to mitigate catastrophic forgetting and mimic the human way of learning [8]. It refers to the learning in which a machine continuously learns from the incoming data. Unlike traditional deep learning, where the complete data is accessible for training at the start, in CL, the data comes sequentially, also known as tasks or episodes. CL allows the acquisition of new knowledge while retaining the existing knowledge. However, there is a tradeoff between how well a model can retain past knowledge and learn new things, called the stability-plasticity dilemma [49]. The plasticity shows the ability of the model to learn new tasks, and the stability refers to maintaining the performance of previously learned tasks upon learning new concepts [50]. Efforts to increase either of them hinder the other and vice versa [9]. More favor for plasticity or stability also depends on the targeted application. There are various CL strategies to mitigate forgetting and allow knowledge accumulation. We discuss them in the following sections. †Section III-A and Section III-B.
III-A Continual learning scenarios
In analysis, particularly audio-based monitoring using anomaly detection, two situations are most prevalent: new samples of the existing class due to various shifts and new intruder/event classes. The first type deals with data with shifted distribution and is termed as ‘new instance’ (NI) scenario; the other type deals with new classes in subsequent episodes and is termed as ‘new class’ (NC) scenario in CL research. Allowing various permutations and combinations on NI, NC, and unique tasks, various CL scenarios are defined across multiple research domains [51], as follows:
III-A1 Data-incremental scenario
Data-incremental scenario aims to continuously learn from new streams of data from an identical underlying distribution during sequential model training. It does not consider new classes or considerable shifts of data distribution and is thus expected to be the easiest among other CL scenarios. It is common for agriculture experts to generate annotations for a dataset in multiple batches, and thus each batch can be considered as a sample of the same data distribution.
III-A2 Domain-incremental scenario
Domain-incremental (DI) scenario deals with situations where the task remains the same, but the context or the information presented changes over time due to the non-stationary environment. In other words, it is an NI scenario with a single task. The data distribution changes, but the classes remain the same for all subsequent episodes. Domain-incremental learning aims to continuously update the model with new streams of data from shifted distributions. Examples of various shift sources in monitoring include variations due to weather, seasons, geo-location, surrounding noise levels across the day, etc.
III-A3 Class-incremental scenario
Class-incremental (CI) scenario is an NC scenario with a single task over the subsequent episodes. It deals with continuously extending the model to unseen classes with new streams of data. Usually, the classes in subsequent episodes are disjoint, and hence the knowledge to learn is also non-overlapping, causing this scenario to be the hardest among CL scenarios. An example of class-incremental learning in real-world monitoring is to identify a cat and dog in one episode and then a cow and horse in the other episode [52].
III-A4 Task-incremental scenario
Task-incremental (TI) scenario refers to learning different kinds of tasks over the episodes. The tasks may also be highly uncorrelated. Each task observed as an episode, usually has a disjoint label space. A real-world example of task-incremental learning is the sequential learning of different sports or musical instruments over time (episode). The information regarding which sports or instrument is the current task can be provided (task-aware) or not required (task-agnostic). The CL literature is ambiguous for task-incremental settings as class-incremental and domain-incremental learning are frequently referred to as task-incremental learning by different research communities [52].
III-A5 Hybrid scenarios
In the real world, it is highly unrealistic that the model will always encounter one type of incremental scenario. It is quite possible that the model can observe new classes as well as domain shifts. The model developed for handling class-incremental situations may not perform well for domain-incremental situations. Thus, there are attempts in the literature to evaluate the same model on different incremental scenarios [53]; however, a separate evaluation is followed for different kinds of incremental learning. In contrast, episodes with a mix of incremental scenarios are closer to real-life situations. Thus it becomes important to also evaluate the CL model in hybrid-incremental settings, i.e., a sequence of episodes with interleaved class-incremental and domain-incremental can be considered for evaluation.
Figure 1: Overview of a continual learning pipeline
III-B Continual learning techniques
Existing CL techniques to handle catastrophic forgetting can be broadly classified into three categories, viz., rehearsal, regularization, and architectural. In rehearsal-based techniques, a small memory is used to store previous episode data and then replayed along with current data to overcome forgetting. The previous data can be stored and replayed in various forms, including raw samples, deep feature descriptors, and generative models to generate samples or features. In contrast, Regularization-based techniques avoid storing past samples and add a regularization term in the loss function or regularize the learning rate to control drastic deviations in learned weights to minimize forgetting. Further, data-focused and prior-focused regularization have been explored. Data-focused approaches use knowledge distillation to transfer past learning to the current model, whereas the prior-focused methods define the importance of the network’s parameters, which is then used to penalize the shift of important parameters [10]. Architectural techniques mainly aim to keep some network parameters reserved for each episode. Two design choices are explored in this category: fixed network capacity and dynamic network capacity. Lastly, hybrid approaches, i.e., a combination of two or more of the above categories, are also popularly explored in various applications [53].
IV CL and non-CL Baselines
The CL research community has contributed various techniques in the above-mentioned categories. Their usefulness has been analyzed in various applications, including object detection [54], robotics [55], medical image analysis [56, 51], etc. For the current application of audio analysis, we have evaluated the following CL and non-CL baselines.
IV-A Continual learning baselines
IV-A1 Elastic Weight Consolidation
Elastic Weight Consolidation (EWC) serves as a regularization technique designed to mitigate the issue of forgetting in neural networks. It achieves this by selectively constraining, or partially freezing, the model weights associated with essential components for previously learned tasks. By prioritizing the preservation of crucial knowledge, EWC helps maintain performance on earlier tasks while accommodating the learning of new ones. This method provides a balance, preventing excessive interference with existing knowledge when adapting to novel information [12, 10].
IV-A2 Learning without Forgetting
Learning without Forgetting (LwF) is a regularization method aimed at addressing the challenge of forgetting in neural networks by enforcing stability in the output or predictions. Notably, LwF prioritizes training efficiency, demonstrating faster results compared to the joint training approach. This is a CL approach, specifically designed to utilize only new data, assuming the unavailability of past data used for initial network pre-training. This sets LwF apart from other CL methods, which typically focus on leveraging prior knowledge to facilitate the learning of new tasks [12][13].
IV-A3 Synaptic Intelligence
Synaptic Intelligence (SI) is treated as a variant of EWC. In contrast to EWC, which estimates parameter importance through a diagonal approximation of the Fisher information, synaptic intelligence takes an online approach, dynamically estimating parameter importance based on the ongoing training trajectory. While EWC relies on a fixed approximation, synaptic intelligence adapts continuously during the learning process, providing a more flexible assessment of parameter significance. This distinction underscores the difference in how these methods evaluate and prioritize the importance of parameters in neural network training [12][11].
IV-A4 Gradient Episodic Memory
Gradient Episodic Memory (GEM) serves as a framework for continual learning, aiming to mitigate catastrophic forgetting by utilizing an episodic memory to retain a portion of the current data samples. By adjusting the gradient of each current task, GEM endeavors to minimize negative backward transfer, thereby preventing the loss of information from previously learned tasks. This approach ensures that the knowledge gained from past tasks positively influences the learning of subsequent tasks, contributing to a more effective and stable learning process [14][57].
IV-A5 Averaged Gradient Episodic Memory
Averaged Gradient Episodic Memory (A-GEM) tries to alleviate the computational cost caused by GEM. A-GEM can be defined as the more efficient version of GEM. Unlike GEM, which minimizes the loss of each previous tasks at each training step by adjusting the gradient of each current task, A-GEM minimizes the average episodic memory loss over the previous tasks at each training step [15].
IV-A6 Greedy sampler and Dumb learner
In Greedy sampler and Dumb learner (GDumb), The sampler greedily stores the part of data samples as they are encountered while ensuring a balanced representation of classes, and the learner which is a neural network trains it dumbly using all the samples available in the memory. Therefore it is called GDumb. Although GDumb is not meant to handle the continual learning scenarios, it gives convincing results in comparison to the CL strategies [16].
IV-A7 Replay
The replay strategy can use different types of replay, like, experience replay where the old stored samples from the memory are used for training with the new samples, generative replay where the old data is generated using some generative models and trained with the new encountered data samples etc. Here we have used simple replay.
IV-B Non-Continual learning baselines
The effectiveness of CL methods is compared against some non-CL approaches that try to mitigate forgetting and achieve high performance using retraining or exploiting all the episode datasets. These approaches do not fit with human brain learning but may perform well at the cost of large amounts of space and computation facilities. We discuss them below.
IV-B1 Joint-training
This is a popular approach in non-CL category, where all the training episodes are jointly trained one time and achieve the highest average performance and hence also regarded as an upper bound. However, this method assumes the availability of all the episodes simultaneously which is a bottleneck.
IV-B2 Cumulative learning
This non-CL training strategy is another widely adopted approach offering upper-bound performance. Here, the training process is started from scratch each time a new episode is available. It considers the current and all the previous episodes together for retraining. However, previously learned knowledge is wasted due to retraining from scratch each time a new episode is encountered leading to high computation requirements.
IV-B3 Naive
This non-CL strategy achieves lower-bound performance. Here, a traditional transfer learning mechanism is followed. Whenever a new episode is available, the model adapts itself to this new data without following any mechanism to control forgetting on past data.
Figure 2: Benchmarking dataset for DI scenario. Here the CL model aims to distinguish between normal and abnormal machine sounds in domain shift conditions.
Figure 3: Benchmarking dataset for CI scenario. Here the classification model aims to distinguish among machine types in class incremental fashion. Here only normal sound of considered machine types is available.
TABLE II: Details of datasets
Continual Learning Scenario Application Tasks #Train clips
#Test clips
Domain-incremental
New
instance
Anomalous
sound detection
T1 (DCASE2021 source),
T2 (DCASE2021 target),
T3 (DCASE2022 source),
T4 (DCASE2022 target),
T5 (DCASE2023 source),
T6 (DCASE2023 target)
3098,
3036,
1512,
1512,
504,
504
862,
844,
420,
420,
140,
140
Class-incremental
New
classes
Machine
sound identification
T1 (ToyCar, ToyConveyor),
T2 (Valve, Fan),
T3 (Pump, slider),
T4 (Vacuum, ToyTank),
T5 (ToyNscale, ToyDrone),
T6 (Bandsaw, Grinder, Shaker)
4320,
4178,
4037,
1425,
1425,
2138
1200,
2361,
3483,
3879,
4275,
4869
V Datasets
To systematically evaluate the CL techniques for monitoring, we require a sequence of datasets to simulate CL scenarios. To the best of our knowledge, no such benchmark datasets have been proposed for audio analysis. Further, there is no large public audio dataset for monitoring; however, there is a well-defined and gigantic repository for abnormality detection in machine sound, namely DCASE. This is very close to the audio-based monitoring application and, hence, suitable for benchmarking CL techniques on audio analysis. Moreover, audio clips comprised various machine sounds in the normal and abnormal categories with and without shifted data distribution due to different machine loads and environmental factors. We thus curate CL benchmark datasets representing different CL scenarios from publically available audio datasets from the DCASE community.
DCASE challenge offers datasets for various audio analysis-related applications every year. Among various tasks, the task-2 of DCASE is for audio anomaly detection and has been held continuously since the year 2020. We have utilized data from the following DCASE challenge tasks from years 2020 to 2023: (a) DCASE2020 task-2: “Unsupervised Detection of Anomalous Sounds for Machine Condition Monitoring”, (b) DCASE2021 task-2: “Unsupervised Anomalous Sound Detection for Machine Condition Monitoring under Domain Shifted Conditions”, (c) DCASE2022 task-2: “Unsupervised Anomalous Sound Detection for Machine Condition Monitoring Applying Domain Generalization Techniques”, and (d) DCASE2023 task-2: “First-Shot Unsupervised Anomalous Sound Detection for Machine Condition Monitoring”. Each of the tasks offers audio clips from 6-7 distinct machine sounds in normal and abnormal classes. The dataset is provided in three segments: development, evaluation, and additional training. We utilized the development segment from each dataset to curate training and testing episodes. Since class-incremental and domain-incremental are the most frequently observed scenarios, therefore we curate these two CL scenarios as described below.
V-A Dataset for DI Scenario
Here, we consider development datasets provided for task-2 of the DCASE2021, DCASE2022, and DCASE2023 challenges as they have the same machine type but have domain shift conditions due to weather and variable machine load. The development dataset has a source domain and then a target domain with a shifted data distribution, which would serve as the domain-incremental condition for our case. Further, the test part of the development segment only offers balanced normal and abnormal classes, and hence, we consider only the test part, not the training part of the development segment. These datasets have normal and abnormal audio clips for each of the 7 machine types, including Fan, Gearbox, Pump, Slider, ToyCar, ToyTrain, and Valve. The end goal here is abnormal sound detection in a sequence of data that may contain a distribution shift over time. We curate six episodes by utilizing source and target domain data from DCASE2021, DCASE2022, and DCASE2023. The first and second episodes represent the source and the target domains of DCASE2021, having 4306 (2153 normal and 2153 abnormal) and 4218 (2109 normal and 2109 abnormal) samples, respectively. Then, the third and fourth episodes are curated from the source and target domain of DCASE2022, offering 2100 (1050 normal and 1050 abnormal) and 2100 (1050 normal and 1050 abnormal) samples, respectively. Lastly, the fifth and sixth episodes are curated from the source and target domain of the DCASE2023 challenge, which has 700 (350 normal and 350 abnormal) and 700 (350 normal and 350 abnormal) samples, respectively. A pictorial representation is provided in Fig. 2 for better visualization of the sequence of episodes under the data-incremental scenario.
V-B Dataset for CI Scenario
Contrary to the domain-incremental dataset discussed in the earlier section, here, the end goal is machine-type classification. In the class-incremental learning scenario, there needs to be the inclusion of classification classes over time, hence we opt to classify machine type as the DCASE datasets offer a large number of machine types which can then be sequentially learned by the CL model. To focus only on machine-type classification, either normal or abnormal data should be considered. Since the normal class data is available in large amounts, hence we consider the normal class samples of different machine sounds here. Specifically, we curate six episodes by utilizing the training part (it has only normal samples) of the development segment in task-2 of DCASE2020 and the source domain of the additional training segment in task-2 of DCASE2023 challenges. The choice of these datasets is such that they have disjoint machine classes. Specifically, we consider ToyCar and ToyConveyer machine types from DCASE2020 in the first episode, Valve and Fan machine types in the second episode, and then Pump and Slider machine types in the third episode. Then from the DCASE2023, two machine types viz., Vacuum and ToyTank are considered in the fourth episode, ToyNscale and ToyDrone are considered in the fifth episode, and finally, the last episode has the remaining classes from DCASE2023, including Bandsaw, Grinder, and Shaker. Inside every episode, the number of samples for each class is considered in balance amount, i.e., the machine types have 3000, 2804, 990, 990, and 990 samples in the first, third, fourth, fifth, and sixth episodes, respectively. Except for the second episode in which the two machine types, namely Valve and Fan, have 3000 and 2804 samples, respectively. A visual representation of the class-incremental dataset is depicted in Fig. 3. The detailed description of datasets of both scenarios is shown in Table II.
VI Experiments
VI-A Experimental setup
In all the approaches, we keep the base model as a pre-trained ResNet50 [58] model initialized with ImageNet weights. We used Adam as an optimizer. In order to select the best hyperparameters for the classification model, multiple experiments were performed with a cumulative approach as it gives the upper bound. Specifically, we search best values for epochs in {10, 15,…50}, batch sizes in {4,8,…32}, and learning rates in {$041e-$ 04, … $051e-$ 05}. Best performance is achieved with a batch size of 8, a learning rate of $041e-$ 04, and epochs of 30 for CI scenario. For DI scenario, the best performance is achieved with a batch size of 8, a learning rate of $031e-$ 03, and epochs of 50. To facilitate a common evaluation scheme the selected hyperparameters are used in all the approaches. Further, the hyperparameters for CL approaches such as $\lambda$, memory size, etc. were also experimentally selected from a range of values. Specifically, for DI experiment, we set $\lambda=0.5$ in EWC, $\alpha=2$, $T=2$} in LwF, and $\lambda=0.8$ in SI. In case of CI experiment, $\lambda=2$ was set in EWC, $\alpha=2$, $T=2$} kept for LwF, and $\lambda=2$ for SI. Further, for both DI and CI, the memory sizes were kept as 2000 samples for GDumb and Replay approaches. Audio clips were trimmed to 10 seconds and then spectrograms were computed on each sample.
Figure 4: The train-test matrix for T=5.
VI-B Evaluation metrics
Here, we describe evaluation procedure and metrics in CL. Each of the training session is a classification problem therefore we compute accuracy on the testing data of each task. After completing the training session of $t^{th}$ task using training data $Tr_{t}$, we compute accuracy values on testing datasets $Te_{1},Te_{2},\dots$ yielding $T$ accuracy values. Thus, after completing all the training sessions on a stream of $T$ tasks, we get a train-test matrix $R^{T\times T}$, where $R_{t,j}$ represents accuracy on $j^{th}$ task after completing the training on $t^{th}$ task. An example of train-test matrix with 5 tasks is shown in Fig. 4.
Apart from getting average accuracy across tasks, other important evaluation measures in CL includes forgetting and forward transfer, which can be computed from this $R$ matrix. Below, we describes metrics for forward transfer, forgetting, and model accuracy.
Backward Transfer (BWT): BWT, a forgetting measure, quantifies the amount of influence of the current task on the previously learned task [14]. It shows the improvement in performance of a task after learning news tasks. Thus, a positive value of BWT signifies that learning of current task $t$ has caused improvement in performance of an old task $k<t$. On the other hand, a negative value of BWT indicates that learning of current task $t$ has caused degradation in performance of an old task. Thus, algorithm with greater values of BWT is superior to the one with small BWT. Blue and yellow cell values in Fig. 4 are utilized to compute BWT using the equation [59] below.
$BWT=\frac{2}{T(T-1)}\sum_{i=2}^{T}\sum_{j=1}^{i-1}(R_{i,j}-R_{j,j})$ (1)
Forward Transfer (FWT): FWT signifies the amount of influence of the current task on upcoming tasks. It shows the zero-shot learning capabilities of the CL model. Higher the value of FWT, better is the model. Red cell values in Fig. 4 are utilized to compute FWT [59] measure with the following equation.
$FWT=\frac{2}{T(T-1)}\sum_{t<j}^{T}{R_{t,j}}$ (2)
Accuracy: Given the train-test matrix $R^{T\times T}$, we compute average accuracy, $ACC$ [14] across tasks after completing the final task $T$ as below:
$ACC=\frac{1}{T}\sum_{j=1}^{T}R_{T,j}$ (3)
Díaz-Rodríguez et. al [59] defines a more rigorous measure of average accuracy by computing $ACC$ after completing each task, termed as $A$. It considers model performances at every training session $t$ and thus better represents the dynamic aspects of CL.
$A=\frac{2}{T(T+1)}\sum_{i\geq j}^{T}R_{i,j}$ (4)
Figure 5: Comparison of different approaches in terms of average accuracy over all episodes after each training session in DI scenario.
TABLE III: Performance comparison for DI scenario. First / second best performance in CL categories indicated in red / blue , respectively. Bold: Upper bound.
Approach
BWT
 [59]
FWT
 [59]
A[59]
ACC
 [14]
CL
EWC ($\lambda=2$)
-5.22 65.09 67.67 71.93
LwF (${\alpha=1,T=1}$)
-1.23 65.85 66.87 68.34
SI ($\lambda=0.5$)
-9.44 63.92 67.31 69.18
GDumb (mem=2000)
-2.25 61.55 61.60 61.42
Replay (mem=2000)
-2.22 68.98 70.12 71.48
GEM (mem=200*6)
-7.40 64.03 66.70 68.43
A-GEM (mem=200*6)
-6.83 64.91 67.28 67.32
Non-CL Naive -9.44 63.91 67.31 69.18
Cumulative -0.17 71.23 72.15 75.40
Joint 76.81
VI-C Results
VI-C1 Performance analysis for DI scenario
For DI scenario, we compute average accuracy over all tasks after completing a training session and report the results via Fig 5. We can see that joint training approach gives the upper bound of performance. CL approaches perform similarly. Further, GDumb strategy and naive training strategy gives lowest performance compared to others. Naive training strategy does not consider any measure to handle forgetting of past tasks and hence lower average performance. On the other hand, GDumb randomly consider a subset data from all encountered tasks for learning from scratch and hence perform poorly due to sample selection bias.
A detailed comparison of best performance achieved by various approaches is provided in Table III. Specifically, for each approach, we report forgetting measure by BWT, forward transfer by FWT, incremental learning capabilities by $A$, and final average accuracy by $ACC$. Here, also we can see that joint and cumulative approaches in non-CL category tend to outperform in all performance metrics as compared to CL approaches. However, assumption of availability of all the tasks together and retraining from scratch whenever a new tasks arrives, limits their applicability in real life. While comparing the CL approaches, we can see that LwF achieves better handles forgetting achieving the best BWT of -1.23. However, other metrics (FWT, $A$, and $ACC$) are lower compared to that achieved by Replay. This is due to the fact that knowledge distillation step in LwF limits the plasticity to learn new tasks. It usually performs well in situation involving less drift/ new data over the episodes (also evident from compromised results in CI scenario). Remarkably, replay outperforms other CL approach including LwF by achieving (FWT, $A$, and $ACC$) as $\{68.98,70.12,71.48\}$. Further, we can see that GDumb with same memory size as with Replay performs poor. GEM and A-GEM approaches are highly compute intensive compared to Replay and also achieve lower performance compared to replay. Although, these CL approaches do not show a contrasting superiority over another for DI scenario. Therefore, CL approaches may be picked based on application requirement, for example, a system demanding a solution with low memory will favour memory-free approaches such as EWC, LwF, or SI and the other way can favour Replay approach as it offers best overall performance.
Figure 6: Comparison of different approaches in terms of average accuracy after each training session in CI scenario.
TABLE IV: Performance Comparison for CI scenario. First / second best performance in CL categories indicated in red / blue , respectively. Bold: Upper bound.
Approach
BWT
 [59]
FWT
 [59]
A[59]
ACC
 [14]
CL
EWC ($\lambda=2$)
-54.60 0.61 10.79 2.07
LwF (${\alpha=2,T=2}$)
-45.72 12.35 20.28 2.96
SI ($\lambda=2$)
-54.77 0 10.13 2.03
GDumb (mem=2000)
-17.58 73.39 77.18 53.15
Replay (mem=2000)
-1.04 96.97 96.98 95.74
GEM (mem=200*6)
-39.87 34.34 42.19 33.01
A-GEM (mem=200*6)
-22.61 62.11 64.82 50.93
Non-CL Naive -54.96 0.55 10.81 2.03
Cumulative -4.93 86.18 85.92 88.22
Joint 90.04
VI-C2 Performance analysis for CI scenario
Figure 6 shows the performance of the different approaches in terms of average accuracy on only seen classes after learning each task in CIS setting. It is expected that over the session performance would drop as more and more classes are added, making the learning difficult. However, Replay is able to maintain the performance to a steady level over the training session. We can see that it outperforms all other CL approaches, even surpassing non-CL strategies supposed to give upper-bound of performance like cumulative and joint.
We report detailed comparison of all approaches in terms of BWT, FWT, $A$, and $ACC$ in Table IV. From the table, we can observe that regularization-based approaches (EWC, LwF, SI) perform similar to the lower-bound approaches i.e., naive approach. We can observe that Replay approach achieves the best values of (BWT, FWT, $A$, $ACC$) as $\{-1.04,96.97,96.98,95.74\}$. It outperforms with an incremental learning accuracy (A) of 96.98% compared to other rehearsal approaches like GDumb, GEM, and A-GEM which achieve only 77.18%, 42.19% and 64.82%, respectively. We can see that GDumb, which is not a carefully designed rehearsal-based approach, acheives second best results and shows superior performance to complex rehearsal-based approaches like GEM and A-GEM.
From tables III and IV, we conclude that the rehearsal-based strategies perform well in both DI and CI scenarios while regularization approaches show inadequate performance in case of CI scenarios. This is attributed to that fact that CI scenario is the most difficult scenario as new task consists of novel classes and hence it is very hard to learn new tasks along with remembering past just with regularization step in the loss function. Some samples from past classes are important to rejuvenate the past class information in the current model.
TABLE V: Performance Comparison with different hyperparameters for DI scenario.
Approach
BWT
 [59]
FWT
 [59]
A[59]
ACC
 [14]
EWC ($\lambda=0.5$) -7.85 64.51 67.03 66.99
EWC ($\lambda=1$) -6.65 63.86 66.79 68.46
EWC ($\lambda=2$) -5.22 65.09 67.67 71.93
LwF (${\alpha=1,T=1}$) -1.23 65.85 66.87 68.34
LwF (${\alpha=2,T=1}$) -1.59 62.65 62.65 63.10
LwF (${\alpha=3,T=1}$) -2.87 60.03 60.97 58.72
SI ($\lambda=0.5$) -9.44 63.92 67.31 69.18
SI ($\lambda=1$) -7.86 63.44 66.50 68.20
SI ($\lambda=2$) -7.86 63.44 66.50 68.20
GDumb (mem=2000) -2.25 61.55 61.60 61.42
GDumb (mem=1500) -0.64 60.21 60.14 58.69
GDumb (mem=1000) -2.36 57.44 57.82 56.62
Replay (mem=2000) -2.22 68.98 70.12 71.48
Replay (mem=1500) -2.40 68.71 69.74 70.80
Replay (mem=1000) -3.65 67.69 69.36 70.31
Replay (mem=500) -3.88 66.69 68.49 70.61
Replay (mem=100) -7.17 64.37 67.02 68.61
Replay (mem=50) -6.67 64.06 66.77 68.00
GEM (mem=10*6) -6.79 63.97 66.68 68.46
GEM (mem=50*6) -5.12 63.93 66.25 69.88
GEM (mem=200*6) -7.40 64.03 66.70 68.43
A-GEM (mem=10*6) -3.72 64.68 66.70 68.99
A-GEM (mem=50*6) -2.66 65.13 67.07 69.12
A-GEM (mem=200*6) -6.83 64.91 67.28 67.32
TABLE VI: Comparison of approaches on different hyperparameters for CI scenario.
Approach
BWT
 [59]
FWT
 [59]
A[59]
ACC
 [14]
EWC ($\lambda=1$) -54.74 0.27 10.44 2.03
EWC ($\lambda=2$) -54.60 0.61 10.79 2.07
EWC ($\lambda=3$) -54.61 0.06 10.17 2.03
LwF (${\alpha=1,T=1}$) -48.68 13.62 21.91 14.31
LwF (${\alpha=2,T=2}$) -45.72 12.35 20.28 2.96
LwF (${\alpha=3,T=3}$) -50.42 9.63 18.39 11.39
SI ($\lambda=1=3$) -54.96 0.55 10.81 2.03
SI ($\lambda=2$) -54.77 0 10.13 2.03
GDumb (mem=2000) -17.58 73.39 77.18 53.15
GDumb (mem=1500) -18.06 71.08 72.92 52.35
GDumb (mem=1000) -12.95 73.59 74.14 52.13
Replay (mem=2000) -1.04 96.97 96.98 95.74
Replay (mem=1500) -1.92 94.75 94.86 93.58
Replay (mem=1000) -1.88 94.50 94.58 93.53
Replay (mem=500) -2.87 94.01 94.30 92.11
GEM (mem=200*6) -39.87 34.34 42.19 33.01
GEM (mem=300*6) -44.82 33.28 41.87 39.00
GEM (mem=400*6) -53.92 21.39 31.76 11.75
A-GEM (mem=200*6) -22.61 62.11 64.82 50.93
A-GEM (mem=300*6) -26.50 59.73 63.34 39.73
A-GEM (mem=400*6) -23.30 60.43 63.29 54.45
VII Discussion
For both DI and CI scenarios, we have conducted various experiments with different techniques, with a wide range of hyperparameters.
From Table V, we observe that different CL strategies perform similarly well in DI scenario. Among the regularization approaches (EWC, LwF, and SI), EWC achieves better performance (BWT, FWT, $A$, $ACC$) as the $\lambda$ value increases, while SI performs better at lower $\lambda$ values. LwF shows an increase in performance (BWT, FWT, $A$, $ACC$) as the $\alpha$ value decreases. EWC outperforms the other regularization approaches in terms of overall accuracy ($A$=67.67%) as well as average accurcay ($ACC$=71.93%), while LwF excels in reducing forgetting (BWT=-1.23) and improving future learning (FWT=65.85).
For rehearsal-based approaches (GDumb, Replay, GEM, and A-GEM), we can see that we can improve performance at the cost of increase memory size. GEM and A-GEM achieve similar performance to regularization approaches (EWC, SI, and LwF) but better than GDumb. Since the compute requirements in GEM and A-GEM are very high, so regularization approaches can be favoured over these. GDumb with lower memory size (mem=1000) show very poor performance and improves if we increase memory size to 2000. It achieves $A$ as 57% and 61.60% with memory size of 1000 and 2000, respectively. In contrast, Replay achieves better results than GDumb even with low memory. It records $A$ as 69.36% and 70.12% for memory size of 1000 and 2000, respectively.
In CI scenario, as shown in Table VI, we observe that regularization-based approaches (EWC, SI, and LwF) fail to mitigate forgetting, for a wide range of hyperparameters as for EWC and SI ($\lambda$ in {1, 2, 3}) and for LwF ($\{\alpha=1,T=1\}$, $\{\alpha=2,T=2\}$, $\{\alpha=3,T=3\}$). Conversely, rehearsal techniques perform relatively well. The performance should increase with increase in memory size but it is sometimes violated in CI scenario by GDumb, GEM and A-GEM. CI scenario is very hard compared to DI and thus a carefully designed strategy succeeds to handle the forgetting. We can observe GDumb perform relatively well compared to GEM and A-GEM. Still, Replay outperforms these approaches by a large margin in terms of all four metrics. On a wide range of memory values, Replay approach consistently outperforms all other techniques used in the study. The best performance in terms of final accuracy $ACC$ by Replay is 95.74% where as that with GDumb, GEM, and A-GEM is 53.15%, 39.00%, and 54.45%, respectively. Further, performance by Replay approach shows a very small decrement even when the memory size is reduced from 2000 samples to 500 samples. A-GEM performs better than GEM in terms of both accuracy and training time when varying the number of samples per task (200, 300, 400). Additionally, the training time for Replay with a memory size of 2000 is significantly less than that of GEM and A-GEM with smaller memory sizes of 1200 (200*6) samples.
Overall, Replay emerges as the most effective technique for CI scenarios, demonstrating consistent performance and efficiency across different memory sizes.
VIII Conclusion
We investigate CL approaches on DI and CI scenarios with 6 sequential tasks curated from a large range of DCASE datasets for machine sound monitoring. With the extensive comparative analysis performed, we observe that the performance of rehearsal-based strategies outperforms regularization-based CL approaches in CI scenario whereas both categories of approaches perform seemingly well in DI scenario. We achieve overall accuracy (A) with Replay approach as 70.12% for the DI scenario and 96.98% for the CI scenario. Overall, this comparative analysis emphasizes the importance of selecting appropriate CL strategies adapted to the specific requirements and complexities of audio data, ensuring optimal performance and robustness in real-world applications.
There are different unexplored real-world challenges associated with the use of CL in audio based application like machine monitoring or farm surveillance. In the future, we aim to investigate CL in farm monitoring which is prone to encounter more challenging and diverse domain shifts due to change is weather, season, and dynamic background environment. Further, in real-life monitoring investigating CL with blurred domain boundaries are yet to be explored.
Acknowledgment
This work is supported by the grant received from DST, Govt. of India for the Technology Innovation Hub at the IIT Ropar in the framework of the National Mission on Interdisciplinary Cyber-Physical Systems.